The sum of the digits of a two digit number is 7. If the digits are reversed, the number is decreased by 27. What is the original number?
Answer by lwsshak3(6519)

The sum of the digits of a two digit number is 7. If the digits are reversed, the number is decreased by 27. What is the original number?
**
let u=units digit
let t=tens digit
original number: 10t+u
reversed number:10u+t
t+u=7
t=7-u
..
10t+u=10u+t-27
10(7-u)+u=10u+7-u-27
70-10u+u=10u+7-u-27
18u=90
u=5
t=7-u=2
original number=10t+u=25